40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 19 18:41:42 2022, yuta, Update, BHD, Contrast measurements for Michelson and ITM-LO ContrastMeasurements.pdf
    Reply  Wed Jul 20 18:07:52 2022, Paco, Update, BHD, BHD MICH test BHD_MICH_OSC.pdf
       Reply  Tue Jul 26 20:54:08 2022, Paco, Update, BHD, BHD MICH test - LO phase control 20220726_BICHD.pdf
          Reply  Wed Jul 27 18:30:50 2022, yuta, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected LOBeamAtBHD.JPGITMXSingleBounceAtBHD.jpg
             Reply  Thu Jul 28 14:34:40 2022, Yehonathan, Update, BHD, LO beam power at BHD DCPDs is significantly lower than expected 
                Reply  Thu Jul 28 16:51:55 2022, Tega, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs Screenshot_2022-07-28_16-58-34_LOandASShaking.pngScreenshot_2022-07-28_18-08-55_DCPDPOPSuspensionCoherence.png
                   Reply  Thu Jul 28 20:16:26 2022, Anchal, Update, BHD, Shaking test for LO beam AS beam to BHD DCPDs 
                   Reply  Fri Jul 29 18:24:53 2022, Tega, Update, BHD, LO beam power improved by factor of 6 after LO and AS beam alignment LO_and_oldAS_settings.pngBHD_fringe_settings.pngScreenshot_2022-07-29_19-37-39.pngFromTheLeft-AS-POP-LO.JPG
                      Reply  Wed Aug 3 16:00:51 2022, yuta, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping Screenshot_2022-08-03_15-46-36_BHDfringeAlmostUnclipped.png
                         Reply  Tue Aug 9 15:33:12 2022, yuta, Update, BHD, BHD fringe contrast improved from 43% to 74% HPC-DCPD_B_OUT_1344118517_ITMY-LO.pngHPC-DCPD_A_OUT_1344118517_ITMY-LO.pngHPC-DCPD_B_OUT_1344118318_ITMX-LO.pngHPC-DCPD_A_OUT_1344118318_ITMX-LO.png
                            Reply  Tue Aug 9 15:50:22 2022, Koji, Update, BHD, BHD fringe contrast improved from 43% to 74% 
                               Reply  Wed Nov 16 15:09:08 2022, yuta, Update, BHD, BHD fringe contrast measured with unwhitening filters ContrastMeasurements20221116_edited.pdf
                         Reply  Tue Oct 25 08:25:00 2022, JC, Update, BHD, BHD fringe aligned with reduced LO and AS beam clipping 
                         Reply  Thu Nov 17 11:24:39 2022, JC, HowTo, LSC, Locking MICH 
                            Reply  Tue Nov 29 11:38:37 2022, JC, HowTo, LSC, Lock Single Arm After MICH lock Screen_Shot_2022-11-29_at_11.01.56_AM.png
Message ID: 17040     Entry time: Wed Jul 27 18:30:50 2022     In reply to: 17037     Reply to this: 17042
Author: yuta 
Type: Update 
Category: BHD 
Subject: LO beam power at BHD DCPDs is significantly lower than expected 

[Paco, Yehonathan, Yuta]

We measured power and counts at BHD DCPDs with LO beam only and ITM single bounce.
We found that LO beam power is ~7 times less than the expected.
We also confirmed that AS beam is clipped somewhere inside vacuum and have 20-50% less power compared with the expected.
LO/AS beams going to DCPD A and B also have power imbalance by 30-40%.

What we did:
 - Run LSCoffsets.py to zero the offsets. I modified the old script so that it can handle new BHD PDs. Also, a bug was fixed (it didn't take into account the gains in filer modules, so INMON is now used instead of OUT16 for calculating offsets).
https://git.ligo.org/40m/scripts/-/blob/main/RFPD/LSCoffsets.py
 - Measured powers and counts in BHD DCPDs at ITMY table with LO beam only and ITMX/ITMY single bounce.
 - During the measurement, we found that power into DCPD A and DCPD B are quite different. One of the reason was a lens and an iris right after the viewport for A path. We removed both of them. Also, only A path have a pickoff which picks off ~20% of light to BHD camera (called SRMF; 40m/16880).
 - We also found that LO beam shape is ugly. ITM single bounce beam from X and Y have similar clipping (see Attached photos). We tried to reduce clipping with various suspensions (LO1, LO2, AS1, AS4, SR2, SRM, BS, ITMX, PR2), but was not possible by moving only single suspension.

Result:
 - Result of counts and power measurements are as follows. Power was measured right in front of DCPD, and also right after the viewport to estimate the loss in the in-air paths. Note that LSC channels have gain of 1, but HPC channels have gain of -1.9 for DCPD_A and -1 for DCPD_B.

                       Blocked  LO       ITMX      ITMY    
C1:LSC-DCPD_A_OUT16    -0.01    -17.89    -91.62    -86.21    
C1:LSC-DCPD_B_OUT16    +0.06    -17.72   -131.83   -131.98    
C1:HPC-DCPD_A_OUT16    +0.07    +34.12   +174.63   +164.24    
C1:HPC-DCPD_B_OUT16    +0.13    +17.60   +131.31   +131.49    
Power at DCPD_A        19 uW    63 uW    278 uW    290 uW    
Power at DCPD_B        19 uW    65 uW    393 uW    404 uW    
Power at viewport A    -- uW    82 uW    350 uW    337 uW    
Power at viewport B    -- uW    64 uW    436 uW    431 uW

DCPD calibration:
 - From the measurements above, counts/W in IN1 can be calculated as follows. Offset of 19 uW is substracted from the measured power to take into account for background light.

C1:LSC-DCPD_A_IN1     -3.59e+05 counts/W
C1:LSC-DCPD_B_IN1     -3.61e+05 counts/W
C1:HPC-DCPD_A_IN1     -3.60e+05 counts/W
C1:HPC-DCPD_B_IN1     -3.57e+05 counts/W

Discussion:
 - DCPD calibration shows that DCPD to ADC itself is quite balanced within 1%. A factor of 1.8-1.9 seen was from unbalanced light between A path and B path (40m/17037).
 - Power expected for ITM single bounce to one of DCPDs is ~520 uW, but was 350-430 uW as measured right after the viewport. Power at A is significantly less than that for B. Note that power at AS55 was as expected (40m/16952). Also, clipping cannot be reduced by moving suspensions. These could mean that clipping is happening after AS2.

950 mW * 0.9 (IMC transmission?) * 5.637%(PRM) * 97.8%(PR2) * 50%(BS) * 98.6%(ITM) * 50%(BS) * 10%(SRM) * 90%(AS2) * 50%(BHDBS) = 520 uW

 - Power expected for LO beam to one of DCPDs is ~530 uW, but was 60-80 uW as measured right after the viewport. Power at A is significantly more than that for B, which is opposite for ITM single bounce. This could mean that something is happening at BHDBS? We are not sure why the power is so low. Are we seeing some ghost beam? For PR2 transmission, 22000 ppm was used for calculation, from 40m/16541.

950 mW * 0.9 (IMC transmission?) * 5.637%(PRM) * 2.2%(PR2) * 50%(BHDBS) = 530 uW

 - As far as we remember, beam shapes were not as bad when we closed out the chambers...

Next:
 - Check if measured power explains the visitivity of LO-ITM single bounce (40m/17020)
 - If not, what is the mode mismatch? Is it possible to explain the mode mismatch with deviations from designed mode-matching telescope?
 - Measure POP power to see if PR2 actually have T=2.2%
 - Play with LO1 and LO2 to invesitate LO beam shape and power
 - Check coherence between LO/AS power fluctuations with suspension motions
 - What is the expected counts/W for these DCPDs?
 - Balance the optical paths in ITMX table for A and B (same lenses, same mirrors)
 - Install better lens in front of camera

Attachment 1: LOBeamAtBHD.JPG  3.896 MB  Uploaded Wed Jul 27 19:57:41 2022  | Hide | Hide all
LOBeamAtBHD.JPG
Attachment 2: ITMXSingleBounceAtBHD.jpg  3.621 MB  Uploaded Wed Jul 27 19:57:52 2022  | Hide | Hide all
ITMXSingleBounceAtBHD.jpg
ELOG V3.1.3-